The perimeter of a square is 72 units. What is the length of the diagonal of the square?
Answer by mananth(16946)

the perimeter of a square =72
4*L =72
L=18
The diagonal and the sides of the square form a right angle triangle.
18^2 + 18^2 = Diagonal^2
diagonal = sqrt(18^2+18^2)
= 25.45 units